# Blind Inverse Problem Solving Made Easy by Text-to-Image Latent Diffusion

### v1

@@ Abstract
Blind inverse problems, where both the target data and forward operator are unknown, are crucial to many computer vision applications. Existing methods often depend on restrictive assumptions such as additional training, operator linearity, or narrow image distributions, thus limiting their generalizability. In this work, we present LADiBI, a training-free framework that uses large-scale text-to-image diffusion models to solve blind inverse problems with minimal assumptions. By leveraging natural language prompts, LADiBI jointly models priors for both the target image and operator, allowing for flexible adaptation across a variety of tasks. Additionally, we propose a novel posterior sampling approach that combines effective operator initialization with iterative refinement, enabling LADiBI to operate without predefined operator forms. Our experiments show that LADiBI is capable of solving a broad range of image restoration tasks, including both linear and nonlinear problems, on diverse target image distributions.
Figure 1: Our proposed LADiBI is a training-free blind inverse problem solving algorithm using large pre-trained text-to-image diffusion models. LADiBI is applicable to a wide variety of image distribution as well as operators with minimal modeling assumptions imposed.
@@ 1 Introduction
Inverse problems are fundamental to many computer vision tasks, where the goal is to recover unknown data ${\bm{x}}$ from observed measurements ${\bm{y}}$. Formally, these problems can be expressed as ${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$, where ${\mathcal{A}}$ is a operator representing the forward process parametrized by $\phi$, and ${\bm{n}}$ is the measurement noise. These problems are critical in fields such as medical imaging, computational photography and signal processing, as they are inherently tied to real-world scenarios like image decompression, deblurring, and super-resolution  [41, 43, 13, 17]. Recently, deep learning-based methods  [30, 22, 44, 8, 7, 10] have emerged as powerful tools for solving inverse problems. In particular, diffusion models offer an elegant solution that incorporates pre-trained models in diverse conditional settings, providing flexible and scalable frameworks for reconstructing high-quality data by guiding the generation process through the measurement constraints [19, 36, 3, 37, 42, 14].
@@
However, most of these efforts focus on inverse problems where the operator ${\mathcal{A}}_{\phi}$ is assumed to be known. In practice, the forward operator is often unknown, leading to what are termed blind inverse problems. Solving these blind inverse problems remains a significant challenge due to their ill-posed nature. Consequently, current methods address this problem by introducing various assumptions about the prior distributions of the target data and unknown operator. These assumptions are typically incorporated through a combination of the following strategies: (1) using explicit formulas or specific neural network architectures to hand-craft inductive biases, (2) simplifying the operator with linear assumptions, (3) restricting target distributions to narrow image classes, (4) collecting measurement, operator or image datasets to train models tailored to specific tasks. While effective in certain cases, these solutions often entail significant computational and manual overhead, limited flexibility, and substantial barriers to practical deployment due to complex model training, costly data curation, unrealistic assumptions, and laborious hyperparameter tuning. These limitations raise a critical question: can we design an algorithm that applies to diverse operators and target image distributions without additional training or data collection?
To tackle this challenge, we formulate this problem as a Bayesian inference problem, where the optimal solution is to sample from the posterior $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi})$. State-of-the-art diffusion-based methods often simplify this problem by assuming independence between the target image and operator distributions [8, 27, 23, 34]. Although this independence assumption reduces the modeling difficulty, we challenge its validity by observing the correlation between the target image distribution and the operator distribution, and argue that this assumption heavily restricted the flexibility of the current diffusion-based approaches. To address this problem effectively, we first propose to estimate the joint distribution $p({\bm{x}},{\mathcal{A}}_{\phi})$ instead. While this estimation seems intractable on the first glance, there actually exists a surprisingly simple solution – large-scale pre-trained text-to-image diffusion models already capture diverse target data and measurement distributions. Moreover, many inverse problems can be described using natural language, allowing tasks like deblurring or super-resolution to be expressed through prompts such as “high-definition, clear image” for the target and “blurry, low-quality” for the measurement. By simply prompting with classifier-free guidance [15], we can approximate the score of this challenging joint prior distributions across a wide range of image and operator types using the same base model, significantly extending the flexibility of the current diffusion-based framework.
In addition to a strong prior, an effective posterior sampling technique is also crucial for meeting measurement constraints. Existing methods generally fall into two categories: Gibbs or EM-style alternating optimization [27, 23], which suits diffusion-based setups without additional training but is sensitive to operator initialization and tuning, and a two-step approach that first estimates the operator followed by solving the non-blind problem [34], which often requires training and strong assumptions. To obtain the best of both worlds, we propose a hybrid approach: first, we obtain a reasonable initial estimate of the operator parameters, and then apply the alternating optimization to iteratively refine both operator parameters and data estimates throughout the diffusion process. Obtaining a reliable initial operator estimate can be difficult for broad functional classes like neural networks. To address this, we introduce a novel initialization scheme that leverages the pseudo-supervision signals from multiple lower-quality data estimation generated by fast posterior diffusion sampling. This approach eliminates the assumption of a specific form of operators, allowing for nonlinear blind inverse problem solving and flexible operator parametrization.
@@
Combining the strong prior with the effective posterior sampling technique, we introduce Language-Assisted Diffusion for Blind Inverse problems (LADiBI), a training-free method that leverages large-scale text-to-image diffusion models to solve a broad range of blind inverse problems with minimal assumptions. LADiBI can be directly applied across diverse data distributions and allows for easy specification of task-specific modeling assumptions through simple prompt adjustments. Algorithm 1 provides an overview of LADiBI, which can be easily adapted from the standard inference algorithm used in popular text-to-image diffusion models. Unlike existing methods, LADiBI requires no model retraining or reselection for different target data distributions or operator functions; instead, all prior parameterization is encoded directly in the prompt, which users can adjust as needed. Notably, we do not assume linearity of the operator, making LADiBI, to the best of our knowledge, the most generalizable approach to blind inverse problem solving in image restoration.
@@
We evaluate LADiBI against state-of-the-art baselines across several blind image restoration tasks, including two linear inverse problems (motion deblurring and Gaussian deblurring) and a nonlinear inverse problem (JPEG decompression), on various image distributions, as illustrated in Figure 1. In the linear cases, our method matches the performance of the state-of-the-art approaches, which require extensive assumptions and model specifications. Additionally, LADiBI is the only method tested that successfully performs JPEG decompression without any information about the task, such as the compression algorithm, the quantization table or factors, relying solely on observations of the compressed images.
@@ 2 Background & Related Works
Method Family Method Prior Type Diverse Image Prior No Additional Training Flexible Operator
Pan-$\ell_{0}$ [28] Inductive bias × ×
Optimization-based Pan-DCP [29] Inductive bias × ×
Unsupervised SelfDeblur [30] Inductive bias × ×
MPRNet [44] Discriminative × ×
Supervised DeblurGANv2 [22] GAN × ×
BlindDPS [8] Pixel diffusion × × ×
GibbsDDRM [27] Pixel diffusion × ×
@@
Diffusion models, also known as score-based generative models [38, 16], generate clean data samples ${\bm{x}}_{0}$ by iteratively refining noisy samples ${\bm{x}}_{t}$ using a time-dependent score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$. This score function is usually parametrized as a noise predictor $\epsilon_{\theta}({\bm{x}}_{t},t)$ and can be used to produce the clean data samples by iteratively denoising the noisy samples from the noise prediction. Particularly, a popular sampling algorithm DDIM [35] adopts the update rule
${\bm{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{{\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{intermediate estimation of ${\bm{x}}_{0}$, denoted as }{\bm{x}}_{0|t}}$
$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{x}}_{t},t)+\sigma_{t}\epsilon$ (1)
@@
Because of this iterative sampling process and its close connection to the score function, many efforts attempt to use unconditionally pretrained diffusion models for conditional generation [38, 26, 11, 37], especially inverse problem solving [9, 7, 19, 36], without additional training. Generally, when attempting to sample from the conditional distribution $p({\bm{x}}|{\bm{y}})$, they decompose its score as
@@
Since $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be obtained from an unconditionally pre-trained diffusio model, these methods usually aim at deriving an accurate approximation for $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$.
As text-to-image latent diffusion models become increasingly popular and powerful  [33, 2, 45], recent work has explored using these large language-conditioned models for inverse problem-solving. In particular, MPGD [14] addresses noisy inverse problems by leveraging the natural manifold preserving property of the latent diffusion models. Specifically, it modifies the intermediate clean latent estimate ${\bm{z}}_{0|t}$ using the following posterior update rule
@@
where ${\mathbf{D}}({\bm{z}}_{0|t})$ represents the decoded intermediate clean image estimation and $c_{t}$ is the step size hyperparameter to account for the unknown noise level $\sigma$. The $L_{2}$ loss $\|{\bm{y}}-{\mathcal{A}}_{\hat{\phi}}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ is induced by the additive Gaussian noise assumption from conventional inverse problem setting. Although these diffusion-based methods performs well on diverse data distributions and tasks, they all require the operator ${\mathcal{A}}_{\hat{\phi}}$ to be known.
@@
where ${\mathcal{A}}_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is the forward operator parameterized by unknown function $\phi$, and ${\bm{n}}\sim{\mathcal{N}}(0,\sigma^{2}I)\in\mathbb{R}^{m}$ represents measurement noise with variance $\sigma^{2}I$. Blind inverse problems are more difficult due to the joint estimation of ${\bm{x}}$ and ${\mathcal{A}}_{\phi}$, and are inherently ill-posed without further assumptions. As a result, existing methods typically incorporate different assumptions about the priors of the target image data as well as the unknown operator.
@@
Conventional methods use hand-crafted functions as image and operator prior constraints [5, 29, 28, 20, 24]. They often obtain these functional constraints by observing certain characteristics (e.g. clear edges and sparsity) unique to certain distributions that are usually considered as the target data distribution (e.g. clear and well-calibrated photos of the real world) and the operator distribution (e.g. some blurring kernels). However, not only are these hand-made functions not generalizable, they also often require significant manual tuning efforts for each individual image.
@@
With the rise of deep learning, neural network has become a popular choice for parameterizing priors [40, 12, 30, 22, 44, 21]. In particular, Kupyn et al. [21, 22], Zamir et al. [44] use supervised learning on data-measurement pairs to jointly learn the image-operator distributions, and Ulyanov et al. [40], Ren et al. [30] apply unsupervised learning solely on measurements by leveraging deep convolutional networks’ inherent image priors. These methods offer significant improvement over traditional approaches. However, their learning procedures require separate data collection and model training for each image distribution and task, making them resource-intensive. Although some attempt to generalize by training on a wide range of data and measurements, our experiments show that they still struggle to adapt to unseen image or operator distributions.
@@
Recently, inspired by advances in diffusion models for inverse problem solving, numerous efforts have incorporated the generative priors from pre-trained diffusion models. However, most of these methods still lack generalizability [8, 27, 34, 23]. For instance, Chung et al. [8], Sanghvi et al. [34] require training separate operator priors, while Murata et al. [27] remains training-free but rely on the operator kernel’s SVD for feasible optimization. These approaches generally depend on linear assumptions about the operator and well-trained diffusion models tailored to specific image distributions, limiting their ability to generalize across diverse image and operator types.
Concurrent to our work, two methods [39, 1] attempt to address these limitations. However, Tu et al. [39] only slightly relaxes the operator constraints to linear function with additive masking. It also only uses pixel diffusion rather than more powerful latent diffusion models. Anonymous [1] relies on hand-crafted update rules for each task, lacking a truly generalizable framework. More importantly, neither method leverages the flexibility provided by text prompts in large text-to-image diffusion models.
@@
Table 1 summarizes the conceptual difference between out method and popular approaches in current literature. By leveraging large-scale pre-trained text-to-image latent diffusion models and our new posterior sampling algorithm, our method offers the most generalizability across diverse image and operator distributions with no additional training.

### v2

@@ Abstract
This paper considers blind inverse image restoration, the task of predicting a target image from a degraded source when the degradation (i.e. the forward operator) is unknown. Existing solutions typically rely on restrictive assumptions such as operator linearity, curated training data or narrow image distributions limiting their practicality. We introduce LADiBI, a training-free method leveraging large-scale text-to-image diffusion to solve diverse blind inverse problems with minimal assumptions. Within a Bayesian framework, LADiBI uses text prompts to jointly encode priors for both target images and operators, unlocking unprecedented flexibility compared to existing methods. Additionally, we propose a novel diffusion posterior sampling algorithm that combines strategic operator initialization with iterative refinement of image and operator parameters, eliminating the need for highly constrained operator forms. Experiments show that LADiBI effectively handles both linear and challenging nonlinear image restoration problems across various image distributions, all without task-specific assumptions or retraining.
@@ 1 Introduction
Figure 1: Our proposed LADiBI is a training-free blind inverse problem solving algorithm for image restoration using large pre-trained text-to-image diffusion models. LADiBI is applicable to a wide variety of image distribution as well as operators with minimal modeling assumptions imposed.
Image restoration is a critical problem in many fields such as medical imaging and computational photography, as it addresses real-world challenges including image decompression, deblurring, and super-resolution  (Yuan et al., 2007; Greenspan, 2009; Isaac and Kulkarni, 2015). These restoration tasks can be formulated as inverse problems, where the goal is to recover unknown image data $\bm{x}$} from observed measurements $\bm{y}$}. Formally, these problems can be expressed as ${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$, where $\mathcal{A}$} is an operator representing the forward degradation process parametrized by $\phi$, and $\bm{n}$} is the measurement noise. Being widely applicable, this problem has attracted numerous solutions, ranging from methods with handcrafted inductive biases to deep learning, especially diffusion-based techniques (Kawar et al., 2022; Song et al., 2022; Bansal et al., 2023; Yu et al., 2023; He et al., 2024).
@@
However, most existing research focuses on the settings where the operator ${\mathcal{A}}_{\phi}$ is known. In practice, the operator is often unknown, leading to what are termed blind inverse problems that present significant challenges due to their ill-posed nature. Current methods attempt to address this problem through several restrictive strategies: (1) introducing hand-crafted inductive biases via explicit formulas or specialized neural architectures (Pan et al., 2018, 2017; Ren et al., 2020), (2) simplifying operators with linear assumptions (Chung et al., 2023a; Murata et al., 2023), (3) constraining target distributions to narrow image classes (Chihaoui et al., 2024; Laroche et al., 2024), or (4) training task-specific models on curated measurement, operator or image datasets (Zamir et al., 2021; Kupyn et al., 2019). While effective in specific scenarios, these approaches suffer from limited flexibility and significant deployment barriers due to costly training and laborious hyperparameter tuning. This raises a fundamental question: can we develop a more generalizable algorithm capable of handling diverse degradation operators and image distributions without additional training or data collection?
To tackle this challenge, we reframe the problem via Bayesian inference, where we sample from the posterior $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi})$. This formulation naturally decomposes the seemingly intractable problem into two simpler sub-problems: estimating the prior $p({\bm{x}},{\mathcal{A}}_{\phi})$ and sampling according to the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$. This decomposition is particularly well-suited for diffusion-based frameworks, where we can leverage pre-trained models as the prior and guide the sampling process via measurement constraints. However, while pre-trained models typically exist for image, adequate priors or training data for the operators are generally unavailable. Thus, existing methods further factorize the joint prior into independent distributions and impose simplifying assumptions such as linearity on the operator, which severely restrict their generality and flexibility.
Instead, our approach is motivated by the key observation that, common restoration tasks can be intuitively described through natural language (e.g., “high-definition, clear image” for targets, “blurry, low-quality” for measurements). Moreover, large pre-trained text-to-image diffusion models already encapsulate rich distributions of both targets and measurements. Although this can, in some way, seem to render the approach less of a true “blind” solution, in practice we note that all methods for blind inverse problems require some assumptions over the space of transformations, and using English text to encode the joint is an extremely flexible and easy-to-use mechanism. Based on these insights, we propose a simple yet powerful method: using classifier-free guidance (Ho and Salimans, 2021), we can approximate the joint prior’s score across diverse images and operators using a single pre-trained text-to-image diffusion model, dramatically broadening the flexibility of existing frameworks. Our approach is particularly advantageous in blind settings, since the appropriate prompts can often be easily inferred directly from the measurements, and textual descriptions enable us to co-encode unknown degradations and desired outputs without training or handcrafted priors.
In addition to estimating the prior score, effectively sampling from the posterior distribution is also crucial for ensuring that restored images satisfy measurement constraints. In blind inverse problems, achieving this requires accurately estimating both data and operator parameters, ideally with generalized parameter classes such as neural networks to improve flexibility. However, reliably initializing these highly unconstrained operator parameters can be challenging. To address this, we propose a novel co-optimization diffusion posterior sampling algorithm specifically tailored for blind inverse problems. Our method begins with a new initialization scheme that leverages pseudo-supervision signals derived from multiple lower-quality target image approximations generated by fast posterior diffusion sampling. We then iteratively refine both operator parameters and data estimates through an alternating optimization procedure integrated within the diffusion sampling process. This strategy eliminates restrictive assumptions about operator forms, thereby enabling nonlinear blind inverse problem solving with highly flexible operator parametrizations.
@@
Combining the text-conditioned prior with our effective posterior sampling, we introduce Language-Assisted Diffusion for Blind Inverse problems (LADiBI), a training-free method that leverages large-scale text-to-image diffusion models to solve a broad range of blind image restoration problems with minimal assumptions. LADiBI can be directly applied across diverse data distributions and allows for easy specification of task-specific assumptions through simple prompting. Algorithm 1 and Figure 2 provide an overview of LADiBI, which can be easily adapted from the standard inference algorithm used in popular text-to-image diffusion models. Unlike existing methods, LADiBI requires no model retraining or reselection for different data distributions or operator functions. Instead, all prior parameterization is encoded directly in the prompt, which users can adjust as needed. Notably, we do not assume linearity of the operator, making LADiBI, to the best of our knowledge, the most generalizable approach to blind inverse problem solving in image restoration.
@@
We evaluate LADiBI against state-of-the-art baselines on a range blind image restoration tasks, including linear problem (e.g. motion and Gaussian deblurring) and nonlinear problem(e.g. JPEG decompression), across various image distributions, as illustrated in Figure 1. In the linear setting, our method matches the performance of the state-of-the-art approaches while requiring significantly fewer assumptions. In the nonlinear setting, LADiBI is the only method tested that can successfully perform JPEG decompression without any prior information of the task (such as the compression algorithm, quantization table or quantization factors), relying solely on observations of the compressed images.
@@ 2 Background & Related Works
Method Family Method Prior Type Diverse Image Prior Training-free Flexible Operator
Pan-$\ell_{0}$ (Pan et al., 2017) Inductive bias × ×
Optimization-based Pan-DCP (Pan et al., 2018) Inductive bias × ×
Self-supervised SelfDeblur (Ren et al., 2020) Inductive bias × ×
MPRNet (Zamir et al., 2021) Discriminative × ×
Supervised DeblurGANv2 (Kupyn et al., 2019) GAN × ×
BlindDPS (Chung et al., 2023a) Pixel diffusion × × ×
BIRD (Chihaoui et al., 2024) Pixel diffusion × ×
GibbsDDRM (Murata et al., 2023) Pixel diffusion × ×
@@
Diffusion models, also known as score-based generative models (Song et al., 2021b; Ho et al., 2020), generate clean data samples ${\bm{x}}_{0}$ by iteratively refining noisy samples ${\bm{x}}_{t}$ using a time-dependent score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$. This score function is usually parametrized as a noise predictor $\epsilon_{\theta}({\bm{x}}_{t},t)$ and can be used to produce the clean data samples through iteratively denoising. Particularly, a popular sampling algorithm DDIM (Song et al., 2021a) adopts the update rule
${\bm{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{{\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{intermediate estimation of ${\bm{x}}_{0}$, denoted as }{\bm{x}}_{0|t}}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{x}}_{t},t)+\sigma_{t}\epsilon$ (1)
@@
Many efforts attempt to use unconditionally pretrained diffusion for conditional generation (Song et al., 2021b; Meng et al., 2022; Dhariwal and Nichol, 2021), especially inverse problem solving (Chung et al., 2023b; Kawar et al., 2022; Song et al., 2022), without additional training. Generally, when sampling from $p({\bm{x}}|{\bm{y}})$, they decompose its score as
@@
Since $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be obtained from an unconditionally pre-trained diffusion model, these methods usually aim at deriving an accurate approximation for $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$.
Recent work has also explored text-to-image latent diffusion models for inverse problem-solving  (Saharia et al., 2022; Balaji et al., 2023; Zhang et al., 2024). In particular, MPGD (He et al., 2024) addresses inverse problems by leveraging the manifold preserving property of the latent diffusion models. Specifically, it modifies the intermediate clean latent estimate ${\bm{z}}_{0|t}$ with
@@
where ${\mathbf{D}}({\bm{z}}_{0|t})$ represents the decoded intermediate clean image estimation and $c_{t}$ is the step size hyperparameter. The $L_{2}$ loss $\|{\bm{y}}-{\mathcal{A}}_{\phi}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ is induced by the additive Gaussian noise assumption from conventional inverse problem setting. Although these diffusion-based methods perform well on diverse data distributions and tasks, they all require the operator ${\mathcal{A}}_{\phi}$ to be known.
@@
where ${\mathcal{A}}_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is the forward degradation operator parameterized by unknown function $\phi$, and ${\bm{n}}\sim{\mathcal{N}}(0,\sigma^{2}I)\in\mathbb{R}^{m}$ represents measurement noise with variance $\sigma^{2}I$. Blind inverse problems are more difficult due to the joint estimation of $\bm{x}$} and ${\mathcal{A}}_{\phi}$, and are inherently ill-posed without further assumptions. As a result, existing methods typically incorporate different assumptions about the priors of the target image data as well as the unknown operator.
@@
Conventional methods use hand-crafted functions as image and operator prior constraints (Chan and Wong, 1998; Pan et al., 2018, 2017; Krishnan et al., 2011; Levin et al., 2009). They often obtain these functional constraints by observing certain characteristics (e.g. clear edges and sparsity) unique to distributions that are usually considered as the target image (e.g. high definition natural images) and the operator (e.g. blurring kernels). However, not only are these hand-made functions not generalizable, they also often require significant manual tuning for each individual image.
@@
With the rise of deep learning, neural network has become a popular choice for parameterizing priors (Ulyanov et al., 2018; Gandelsman et al., 2019; Ren et al., 2020; Kupyn et al., 2019; Zamir et al., 2021). In particular, Kupyn et al. (2019); Zamir et al. (2021) use supervised learning on data-measurement pairs to jointly learn the image-operator distributions, and Ulyanov et al. (2018); Ren et al. (2020) apply self-supervised learning solely on measurements using deep convolutional networks’ inherent image priors. These methods offer significant improvement over traditional approaches. However, their learning procedures require separate data collection and model training for each image distribution and task, making them resource-intensive. Although some attempt to generalize by training on a wide range of data and measurements, our experiments show that they still struggle to adapt to unseen image or operator distributions.
@@
Recently, inspired by advances in diffusion models for inverse problem solving, numerous efforts have incorporated the priors from pre-trained diffusion models. However, most of these methods still lack generalizability (Chung et al., 2023a; Murata et al., 2023; Sanghvi et al., 2023; Laroche et al., 2024; Tu et al., 2024; WeiminBai et al., 2025). For instance, Chung et al. (2023a); Sanghvi et al. (2023) require training separate operator priors, while Murata et al. (2023) remains training-free but rely on the operator kernel’s SVD for feasible optimization. These approaches generally depend on linear assumptions about the operator and well-trained diffusion models tailored to specific image distributions, limiting their ability to generalize across diverse image and operator types.
@@
Table 1 summarizes the conceptual difference between our method and popular approaches in current literature. By leveraging large-scale pre-trained text-to-image latent diffusion models and our new posterior sampling algorithm, our method offers the most generalizability across diverse image and operator distributions with no additional training.